OpenStudy (anonymous):

find the point of intersection 3f+2g=30 and f+2g=26

OpenStudy (anonymous):

first solve for g Then you can easily put it in your calculator :]

OpenStudy (anonymous):

or is that the part you need help on?

OpenStudy (anonymous):

yes. please

OpenStudy (anonymous):

sorry my computer doesn't like the drawing apperently! Okay,

OpenStudy (anonymous):

im going to make g = y and f = x okay?

OpenStudy (anonymous):

3x + 2y = 30 2y = 3x +30 y= (2/3)x + 15

OpenStudy (anonymous):

and x+ 2y = 26 2y = -x + 26 y = (-1/2)x + 13

OpenStudy (anonymous):

then you put both in the calculator. Hit graph, 2nd, Calc, and intersect :]

OpenStudy (anonymous):

wow thanks so much, one quick ques.. for the first equation, i believe u divided by 2 correct? (3x+2y=30). so then shouldnt it be 3/2x +15?